OpenStudy (justtryingtosurvive):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for three different years: Number of years 1 2 3 Option 1 (amount in dollars) 1300 1690 2197 Option 2 (amount in dollars) 1300 1600 1900 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years.

OpenStudy (hanna1):

Part A- Option 1 is exponential and increases 130% per year. Option 2 increases $300 per year and it's linear. Part B- f(n)=1000+300n option 2 f(n)=1000(1.30)^n option 1